(5-y)^2+y^2=13
(5-y)*(5-y)+y*y=13
25-5y-5y+y^2+y^2=13
25-10y+2y^2=13
2y^2-10y=13-25
2y^2-10y=-12
2y^2-10y+12=0
factorizing;
y=2 or 3

Found 3 solutions by Queen Dorex, JBarnum, almiskinbillah:
Answer by Queen Dorex(2)

(5-y)^2+y^2=13
(5-y)(5-y) + y^2 =13
25-10y+2y^2=13
2y^2-10y+12=0
y=2 or 3

Answer by JBarnum(2146)

(5-y)(5-y)+y^2=13
Foil method
25-10y+y^2 + y^2=13
Subtract 13 and add like terms
2y^2-10y+12=0
You can solve it this way with quadratic formula if you want but I'm gonna do it the faster easy way
Factor out a 2
2(y^2-5y+6)=0
Find two numbers that multiply to get +6 and add to get -5
-3 and -2
2(y-3)(y-2)=0
y=3 , 2

Answer by almiskinbillah(1)